# Co-evolution of Nuclear Star Clusters and Massive Black Holes: Extreme Mass-Ratio Inspirals

### v1

@@ 3.2.4 Mass growth of the MBH
The mass increase of the MBH at $12$ Gyr is approximately the sum of $\delta M_{\bullet,\rm td}$ (mass contributed by accreting gas from TDEs) and $\delta M_{\bullet,\rm ml}$ (mass contributed by stellar mass loss). Then the final mass of the MBH at $12$ Gyr is
$M_{\bullet,f}$ $\simeq\delta M_{\bullet,\rm td}+\delta M_{\bullet,\rm ml}+M_{\bullet,i}$ (10)
$\simeq\langle m_{\star}\rangle_{\rm td}\bar{R}_{\rm td,MS}T_{12\rm\,Gyr}+M_{\rm cl,i}f_{\rm ml}f_{\rm ma}+M_{\bullet,i}$
@@
  • $\langle m_{\star}\rangle_{\rm td}\sim 0.5M_{\odot}$ is the mean mass of TDEs (see Section 3.2.3);
@@
  • $T_{12\rm\,Gyr}=12$ Gyr is the time of accretion;
@@
  • $f_{\rm ml}=0.4\sim 0.7$ (see Section 3.2.2) is the fraction of stellar mass lost relative to the initial cluster mass;
@@
According to Section 3.2.3 and Figure 8, the mean MS-TDE rates are generally $\bar{R}_{\rm td,MS}\sim 10^{-3}$ yr-1, $\sim 10^{-4}$ yr-1, and $\sim 10^{-6}$ yr-1  for massive NSCs ($M_{\rm cl,i}=10^{9}M_{\odot}$), Milky-Way-like NSCs, and smaller NSCs ($M_{\rm cl,i}=2\times 10^{6}M_{\odot}$), respectively. The total mass contributed by TDEs over $12$ Gyr is typically $\delta M_{\bullet,\rm td}\sim 10^{7}M_{\odot}$, $\sim 10^{6}M_{\odot}$, and $\sim 5\times 10^{4}M_{\odot}$ for massive, Milky-Way-like, and smaller NSCs, respectively. These results are consistent with the final MBH mass for models with $f_{\rm ma}=0$, e.g., models MKG91F0k, MKG82F0k, and MKG71F0k.
@@ 3.2.5 Reproducing the Milky Way’s NSC
As discussed in Section 3.2.4, the mass growth of the MBH is contributed mainly by both TDEs of MSs (which can be regulated by the initial effective size of the cluster, $r_{\rm eff,i}$) and mass loss from stellar evolution (which depends critically on the assumed value of $f_{\rm ma}$). Model MKG81F0k ignores stellar mass loss ($f_{\rm ma}=0$) and thus requires a compact initial cluster size ($r_{\rm eff,i}=1.6$ pc) to reproduce the observed MBH mass at $12$ Gyr solely via TDEs. Model MKG82F0.06k includes stellar mass loss with an assumed value of $f_{\rm ma}=0.06$ and thus can initially have a slightly larger cluster size.
@@
3.2.6 The evolution of EMRI event rates
@@
  • The size expansion or contraction of the cluster (see also Section 3.2.2). The expansion of the cluster’s size due to either substantial stellar mass loss or relaxation processes leads to a decrease in the cluster’s density. As $R_{\rm EMRI}\propto M_{\rm cl}/r_{\rm eff}^{3}$ approximately, the EMRI rates for all stellar objects are expected to decline. However, if the growth of the MBH mass is sufficiently fast such that the local relaxation time is longer than the $e$-folding time of the MBH’s mass ($\sim 4.5\times 10^{7}$ yr), the shrinking of stellar orbits can slow down or even reverse the size expansion. If the latter occurs, the EMRI event rates can increase. This effect is particularly apparent in model MKG83F1k, where all EMRI rates show a very substantial increase around $\sim 0.5$ Gyr.
  • The evolution of the stellar population (see also Section 3.2.1). Almost all SBHs and NSs form in the early universe, and afterward, the sizes of their populations remain almost constant over cosmic time. Thus, their rates are affected mainly by the dynamical evolution of the cluster and the MBH. However, the rate evolution of WD-EMRIs is slightly complicated by the growing number of their populations over cosmic time (See Figure 6).
  • The IMF of the cluster. The IMF affects the evolution of the number and mass distribution of stellar populations over cosmic time (see Section 3.2.1). Different IMFs lead to different amounts of stellar mass loss, which affects the size expansion of the cluster and the mass growth of the MBH. These complications ultimately lead to differences in the observed rates and mass distribution of EMRIs.
@@
  • For BDs and MSs, they become EMRIs only if $M_{\bullet}\gtrsim 10^{5}M_{\odot}$ (see Section 3.1.2). The ISO pericenters of BDs and MSs are outside their tidal radii only if $M_{\bullet}\gtrsim 10^{6}M_{\odot}$ (see Figure 1). Around smaller MBHs, they are more likely to be tidally disrupted when passing through the loss cone. Thus, starting from a seed MBH of $10^{4}M_{\odot}$, BD-EMRIs and MS-EMRIs appear at later epochs of the Universe. BD-EMRIs and MS-EMRIs form more easily around more massive MBHs, as their tidal radii become relatively smaller compared to the ISO. However, how their rates evolve with time is also affected by the factors listed above.
@@
The rates of SBH-EMRIs are orders of magnitude larger than those of other stellar types, which can be understood from the analysis of the critical radius $a_{\rm crit}$ discussed in Section 3.1.2.
@@
Amaro-Seoane (2019) predicted a merger rate of $\dot{\Gamma}_{\rm X-MRI}\sim 10^{-6}\,\textrm{yr}^{-1}$ for the Galactic Center, assuming $\beta=1.5$ for BDs and $\gamma=1.75$ for SBHs. Our simulations for Milky-Way-like models (e.g., MKG82F0.06k) yield BD-EMRI rates of $\sim 10^{-9}$ yr-1 at 12 Gyr. This discrepancy arises for three reasons: (1) Amaro-Seoane (2019) assumes a steady solution and no MBH mass growth, so it is more appropriate to compare the results with the initial rates of model M5_2 in Section 3.1.3. As shown in the bottom panel of Figure 4, the initial BD-EMRI rate is around $\sim 10^{-7}$ yr-1, consistent with the results of Amaro-Seoane (2019) within one order of magnitude. (2) In GNC, the rates are time-dependent, and are mainly affected by the cluster expansion, the mass growth of MBH and the assumed IMF (see also ection 2.6). The expansion of the cluster reduces the central density of BDs, leading to lower rates of EMRIs. (3) By $12$ Gyr, the density slope of BDs in GNC is $\beta\sim 1.3$ (See right panel of Figure 10), which is flatter than the assumed slope of $1.5$. A flatter density profile results in smaller rates of EMRIs (See Equation 9).
However, the central argument of Amaro-Seoane (2019) concerns the steady-state population observable by LISA. Given the extremely long residence time of X-MRIs (up to $2\times 10^{6}$ years in the LISA band), they calculated a population of $\sim 20$ observable sources in the Galactic Center, assuming the system is currently relaxed. While the long-term evolution in our simulations leads to lower present-day merger rates than the idealized analytical model, the dynamical pathway for forming X-MRIs is validated. The analytical estimate of the steady-state population remains a relevant prediction for the current epoch—the timeframe when LISA will operate—provided the relaxation processes in the Galactic Center today are efficient enough to support the required influx, as the conditions will not change significantly over the residence timescale.
@@
3.2.7 The mass and eccentricity distribution of EMRIs in the LISA band
@@
Here we investigate the present-day (at $12$ Gyr in the simulation) mass and orbital eccentricity distributions of EMRIs in the LISA band. We focus on the eccentricity of EMRIs ($e_{\rm 0.1mHz}$) when their orbital frequencies have decayed to $f_{\rm orb}=0.1$ mHz ($P=10^{4}$ s). Their primary GW emissions are then around $0.2$ mHz, which is within LISA’s sensitivity range.
The characteristic orbital semi-major axis of these EMRIs is
$a_{\rm 0.1mHz}\sim 18.45~{\rm rg}\times\left(\frac{P}{10^{4}~{\rm s}}\right)^{1/3}\left(\frac{M_{\bullet}}{4\times 10^{6}~M_{\odot}}\right)^{-2/3}.$ (11)
@@
BDs are more likely to become EMRIs if they are more massive. For less massive BDs (e.g., $m\lesssim 0.05M_{\odot}$), their physical sizes are larger, and their GW orbital decay is so weak that the critical distances $a_{\rm crit}$ defined in Section 3.1.2 are much smaller. Below the maximum eccentricity $e_{\rm 0.1mHz,max}$, their eccentricity distribution peaks at lower eccentricities compared to compact objects.
@@
3.2.8 The inclination distribution of EMRIs in the LISA band
@@
4 Discussion
By comparing the model results from Section 3.1 and the results presented in Section 4 of Paper II with those in Section 3.2, we can see that ignoring stellar evolution leads to significant underestimates of the cluster’s size expansion, the loss of the NSC’s total mass, and the mass growth of the MBH. Thus, our results demonstrate that stellar evolution must be included to achieve realistic simulations of NSC dynamics and MBH evolution, which is crucial for accurate predictions of TDE and EMRI properties.
@@ 4 Discussion
However, the analytical models provide valuable estimates for the currently observable population, which is the relevant timeframe for the upcoming LISA mission. Provided the nuclear star cluster is currently relaxed, the steady-state analysis offers a snapshot of the expected EMRI population. The GNC simulations confirm the underlying dynamical mechanisms assumed in the analytical work. The analytical estimates of the steady-state population (Amaro Seoane et al. (2024), Amaro-Seoane (2019)) remain valid estimates (within an order of magnitude) for the present epoch, as the astrophysical conditions are not expected to change significantly over the relatively short residence timescales of the EMRIs (e.g., the next $10^{5}$ years), even though the simulations demonstrate that these conditions evolve over a Hubble time.
5 Conclusion
@@
6 Acknowledgments
We thank Matteo Sadun Bordoni and Sebastiano Fellenberg for helpful discussions of this work. This work was supported in part by the National Natural Science Foundation of China under grant Nos. 12273006 and 11833007. This work was also supported in part by the Key Project of the National Natural Science Foundation of China under grant No. 12133004. The simulations in this work were performed partly at the TianHe-II National Supercomputer Center in Guangzhou.
Appendix A Solving for the Relativistic ISO
Figure A: Left panel: The pericenter distance $r_{p}$ (in units of $r_{g}$) of an innermost stable orbit (ISO) in Boyer-Lindquist (BL) coordinates, plotted as a function of the orbital semi-major axis, defined as $a_{\rm GR}=(r_{p}+r_{a})/2$, where $r_{a}$ is the apocenter distance. Both $r_{p}$ and $r_{a}$ are solved using the method described in Appendix A. The colored lines show the results for the case of $a=1$ with various inclination angles ii. The solid black line shows the result for a Schwarzschild MBH. Middle panel: The mapped Newtonian pericenter distance $r_{\rm p,NW}$ of the ISO as a function of the Newtonian semi-major axis $a_{\rm NW}$, both defined by Equation B1. Right panel: The Newtonian pericenter distance $r_{\rm p,NW}$ of the ISO as a function of the Newtonian inclination angle $i_{\rm NW}$, calculated for orbits with a semi-major axis of $a_{\rm NW}=10^{4}r_{g}$.

### v2

@@ III.2.4 Mass growth of the MBH
The mass increase of the MBH at $12$ Gyr is approximately the sum of $\delta M_{\bullet,\rm td}$ (mass contributed by accreting gas from TDEs), $\delta M_{\bullet,\rm d}$ (mass contributed by direct swallowing of stellar objects) and $\delta M_{\bullet,\rm ml}$ (mass contributed by stellar mass loss). Then the final mass of the MBH at $12$ Gyr is
$M_{\bullet,f}\simeq\delta M_{\bullet,\rm td}+\delta M_{\bullet,\rm d}+\delta M_{\bullet,\rm ml}+M_{\bullet,i}$ (8)
$\simeq$ $\langle m_{\star}\rangle_{\rm td}\bar{R}_{\rm td,MS}T_{12\rm\,Gyr}+\langle m\rangle_{\rm d}\bar{R}_{\rm d}T_{12\rm\,Gyr}$
$+$ $M_{\rm cl,i}f_{\rm ml}f_{\rm ma}+M_{\bullet,i}$
@@
  • $\langle m_{\star}\rangle_{\rm td}\sim 0.5M_{\odot}$ is the mean mass of TDEs (see Section III.2.3);
@@
  • $T_{12\rm\,Gyr}=12$ Gyr is the time of evolution;
  • $\langle m\rangle_{\rm d}$ and $\bar{R}_{\rm d}$ is the mean mass and event rate of objects directly swallowed by the central MBH;
@@
  • $f_{\rm ml}=0.4\sim 0.7$ (see Section III.2.2) is the fraction of stellar mass lost relative to the initial cluster mass;
@@
According to Section III.2.3 and Figure 8, the mean MS-TDE rates are generally $\bar{R}_{\rm td,MS}\sim 10^{-3}$ yr-1, $\sim 10^{-4}$ yr-1, and $\sim 10^{-6}$ yr-1  for massive NSCs ($M_{\rm cl,i}=10^{9}M_{\odot}$), Milky-Way-like NSCs, and smaller NSCs ($M_{\rm cl,i}=2\times 10^{6}M_{\odot}$), respectively. The total mass contributed by TDEs over $12$ Gyr is typically $\delta M_{\bullet,\rm td}\sim 10^{7}M_{\odot}$, $\sim 10^{6}M_{\odot}$, and $\sim 5\times 10^{4}M_{\odot}$ for massive, Milky-Way-like, and smaller NSCs, respectively. These results are consistent with the final MBH mass for models with $f_{\rm ma}=0$, e.g., models MKG91F0k, MKG82F0k, and MKG71F0k.
@@
The total rate of objects falling into the loss cone is given by $\bar{R}_{\rm lc}\simeq\bar{R}_{\rm td,MS}+\bar{R}_{\rm d}$. The contribution from the direct swallowing of objects is usually smaller than that from TDEs. However, it can become significant and even larger than those from TDEs for MBHs with masses $M_{\bullet}\gtrsim 10^{7}M_{\odot}$. For example, in model MKG91F0k, which has a final MBH’s mass of $M_{\bullet}\sim 2\times 10^{7}M_{\odot}$, the contribution from direct capture can reach up to $\sim 60\%$.
@@ III.2.5 Reproducing the Milky Way’s NSC
As discussed in Section III.2.4, the mass growth of the MBH is contributed mainly by both TDEs of MSs (which can be regulated by the initial effective size of the cluster, $r_{\rm eff,i}$) and mass loss from stellar evolution (which depends critically on the assumed value of $f_{\rm ma}$). Model MKG81F0k ignores stellar mass loss ($f_{\rm ma}=0$) and thus requires a compact initial cluster size ($r_{\rm eff,i}=1.6$ pc) to reproduce the observed MBH mass at $12$ Gyr solely via TDEs. Model MKG82F0.06k includes stellar mass loss with an assumed value of $f_{\rm ma}=0.06$ and thus can initially have a slightly larger cluster size.
@@
III.2.6 The evolution of EMRI event rates
@@
  • The size expansion or contraction of the cluster (see also Section III.2.2). The expansion of the cluster’s size due to either substantial stellar mass loss or relaxation processes leads to a decrease in the cluster’s density. As $R_{\rm EMRI}\propto M_{\rm cl}/r_{\rm eff}^{3}$ approximately, the EMRI rates for all stellar objects are expected to decline. However, if the growth of the MBH mass is sufficiently fast such that the local relaxation time is longer than the $e$-folding time of the MBH’s mass ($\sim 4.5\times 10^{7}$ yr), the shrinking of stellar orbits can slow down or even reverse the size expansion. If the latter occurs, the EMRI event rates can increase. This effect is particularly apparent in model MKG83F1k, where all EMRI rates show a very substantial increase around $\sim 0.5$ Gyr.
  • The evolution of the stellar population (see also Section III.2.1). Almost all SBHs and NSs form in the early universe, and afterward, the sizes of their populations remain almost constant over cosmic time. Thus, their rates are affected mainly by the dynamical evolution of the cluster and the MBH. However, the rate evolution of WD-EMRIs is slightly complicated by the growing number of their populations over cosmic time (See Figure 6).
  • The IMF of the cluster. The IMF affects the evolution of the number and mass distribution of stellar populations over cosmic time (see Section III.2.1). Different IMFs lead to different amounts of stellar mass loss, which affects the size expansion of the cluster and the mass growth of the MBH. These complications ultimately lead to differences in the observed rates and mass distribution of EMRIs.
@@
  • For BDs and MSs, they become EMRIs only if $M_{\bullet}\gtrsim 10^{5}M_{\odot}$ (see Section III.1.2). The ISO pericenters of BDs and MSs are outside their tidal radii only if $M_{\bullet}\gtrsim 10^{6}M_{\odot}$ (see Figure 1). Around smaller MBHs, they are more likely to be tidally disrupted when passing through the loss cone. Thus, starting from a seed MBH of $10^{4}M_{\odot}$, BD-EMRIs and MS-EMRIs appear at later epochs of the Universe. BD-EMRIs and MS-EMRIs form more easily around more massive MBHs, as their tidal radii become relatively smaller compared to the ISO. However, how their rates evolve with time is also affected by the factors listed above.
@@
The rates of SBH-EMRIs are orders of magnitude larger than those of other stellar types, which can be understood from the analysis of the critical radius $a_{\rm crit}$ discussed in Section III.1.2.
@@
Amaro-Seoane (2019) predicted a merger rate of $\dot{\Gamma}_{\rm X-MRI}\sim 10^{-6}\,\textrm{yr}^{-1}$ for the Galactic Center, assuming $\beta=1.5$ for BDs and $\gamma=1.75$ for SBHs. Our simulations for Milky-Way-like models (e.g., MKG82F0.06k) yield BD-EMRI rates of $\sim 10^{-9}$ yr-1 at 12 Gyr. This discrepancy arises for three reasons: (1) Amaro-Seoane (2019) assumes a steady solution and no MBH mass growth, so it is more appropriate to compare the results with the initial rates of model M5_2 in Section III.1.3. As shown in the bottom panel of Figure 4, the initial BD-EMRI rate is around $\sim 10^{-7}$ yr-1, consistent with the results of Amaro-Seoane (2019) within one order of magnitude. (2) In GNC, the rates are time-dependent, and are mainly affected by the cluster expansion, the mass growth of MBH and the assumed IMF (see also ection 2.6). The expansion of the cluster reduces the central density of BDs, leading to lower rates of EMRIs. (3) By $12$ Gyr, the density slope of BDs in GNC is $\beta\sim 1.3$ (See right panel of Figure 10), which is flatter than the assumed slope of $1.5$. A flatter density profile results in smaller rates of EMRIs (See Equation 7).
Thus, the discrepancies are mainly ascribed to the differences between the cluster conditions assumed in the steady-state solution and those obtained from our numerical simulation. Nevertheless, our numerical simulations validate the dynamical pathway for forming X-MRIs. Considering that our simulation results are still subject to several uncertainties, e.g., in the initial cluster’s conditions and in other complexities not captured in our simulation (see Section IV), we do not intend to exclude these steady-state solutions.
@@
III.2.7 The mass and eccentricity distribution of EMRIs in the LISA band
@@
Here we investigate the present-day (at $12$ Gyr in the simulation) mass and orbital eccentricity distributions of EMRIs in the LISA band. We focus on the eccentricity of EMRIs ($e_{\rm 0.1mHz}$) when their orbital frequencies have decayed to $f_{\rm orb}=0.1$ mHz ($P=10^{4}$ s). Their primary GW emissions are then $\sim 0.2$ mHz for circular orbits (or $\sim 5$mHz if $e=0.9)$, which is within LISA’s sensitivity range.
The orbital semi-major axis of these EMRIs for a given period ($P$) is
$a_{P}\sim 18.45~{r_{g}}\times\left(\frac{P}{10^{4}~{\rm s}}\right)^{2/3}\left(\frac{M_{\bullet}}{4\times 10^{6}~M_{\odot}}\right)^{-2/3}.$ (9)
@@
BDs are more likely to become EMRIs if they are more massive. For less massive BDs (e.g., $m\lesssim 0.05M_{\odot}$), their physical sizes are larger, and their GW orbital decay is so weak that the critical distances $a_{\rm crit}$ defined in Section III.1.2 are much smaller. Below the maximum eccentricity $e_{\rm 0.1mHz,max}$, their eccentricity distribution peaks at lower eccentricities compared to compact objects.
@@
Note that the eccentricities discussed here are defined within Newtonian physics (see more details in Appendix C). A more accurate definition for eccentricity under relativistic framework should be given by the pericenter and apocenter of the relativistic orbit, e.g., $e=(r_{a}-r_{p})/(r_{a}+r_{p})$, where $r_{a}$ and $r_{p}$ are the two roots of Equation B1. Usually the relativistic eccentricities are larger than those of Newtonian ones, as the pericenter can be closer than those in the Newtonian. Nevertheless, we still observe similar trends of bias towards high-eccentricity sources in the eccentricity distribution, which has been observed in some studies that adopt relativistic orbits and a Schwarzschild MBH (e.g., Hopman & Alexander, 2005; Mancieri et al., 2025). We defer a full relativity study of the eccentricity and inclination distributions of EMRIs around non-spinning and spinning MBH to future works.
III.2.8 The inclination distribution of EMRIs in the LISA band
@@
IV Discussion
By comparing the model results from Section III.1 and the results presented in Section 4 of Paper II with those in Section III.2, we can see that ignoring stellar evolution leads to significant underestimates of the cluster’s size expansion, the loss of the NSC’s total mass, and the mass growth of the MBH. Thus, our results demonstrate that stellar evolution must be included to achieve realistic simulations of NSC dynamics and MBH evolution, which is crucial for accurate predictions of TDE and EMRI properties.
@@ IV Discussion
However, the analytical models provide valuable estimates for the currently observable population, which is the relevant timeframe for the upcoming LISA mission. The GNC simulations confirm the underlying dynamical mechanisms assumed in the analytical work. Provided some suiable conditions of the nuclear star cluster (such as the model M5_2 in Figure 4), the analytical estimates of the steady-state population (Amaro Seoane et al. (2024), Amaro-Seoane (2019)) remain valid estimates (within an order of magnitude) for the present epoch. These steady-state analysis offers a snapshot of the expected EMRI population. as the astrophysical conditions are not expected to change significantly over the relatively short residence timescales of the EMRIs (e.g., the next $10^{5}$ years), even though the simulations demonstrate that these conditions evolve over a Hubble time.
V Conclusion
@@
VI Acknowledgments
We thank Matteo Sadun Bordoni and Sebastiano Fellenberg for helpful discussions of this work. This work was supported in part by the National Natural Science Foundation of China under grant Nos. 12273006. This work was also supported in part by the Key Project of the National Natural Science Foundation of China under grant No. 12133004. The simulations in this work were performed partly at the TianHe-II National Supercomputer Center in Guangzhou.
Appendix A Numerical realization of gravitational wave orbital decay
A.1 For orbits bound to the MBH
For stellar orbits bound to the MBH, the decay of energy and angular momentum per orbit, $\overline{\delta E}^{\rm GW}$ and $\overline{\delta J}^{\rm GW}$, are given by (Peters & Mathews, 1963; Peters, 1964):
$\overline{\delta E}^{\rm GW}=D^{\rm GW}_{E}P=-\frac{64\pi}{5}\frac{M_{\bullet}^{2}m(m+M_{\bullet})^{1/2}G^{7/2}}{r_{p}^{7/2}c^{5}}\frac{1+\frac{73}{24}e^{2}_{K}+\frac{37}{96}e^{4}_{K}}{(1+e_{K})^{7/2}},$ (A1)
$\overline{\delta J}^{\rm GW}=D^{\rm GW}_{J}P=-\frac{64\pi}{5}\frac{M_{\bullet}^{2}mG^{3}}{r_{p}^{2}c^{5}}\frac{1+\frac{7}{8}e^{2}_{K}}{(1+e_{K})^{2}},$
This can be realized in our numerical simulations as follows: If $j>0.4$ ($e\lesssim 0.9$), we set
$\delta E^{\rm GW}=D^{\rm GW}_{E}\delta t,$ $\delta J^{\rm GW}=D^{\rm GW}_{J}\delta t$ (A2)
If $j<0.4$ ($e\gtrsim 0.9$), we set
$\delta E^{\rm GW}=D^{\rm GW}_{E}Pn_{p},$ $\delta J^{\rm GW}=D^{\rm GW}_{J}Pn_{p}$ (A3)
where
$n_{p}$ $=\left\{\begin{array}[]{cc}k+1&{\rm if~}\mathcal{M}_{i}<0~~{\rm and}~~\mathcal{M}_{i}+\delta\mathcal{M}\geq 0\\ k&{\rm otherwise}\\ \end{array}\right.,~~~\mathcal{M}_{f}=\mathcal{M}_{i}+2\pi k+\delta\mathcal{M}$ (A4)
where $\mathcal{M}_{i}$ and $\mathcal{M}_{f}$ are the initial and final mean anomaly of the orbit during a timestep $\delta t$, respectively.
To ensure convergence we additionally place the following two constraints on the particle’s time step $\delta t$ (in addition to Equation B1-B3 of Zhang & Amaro Seoane (2024)), such that
[…]
where
$\delta t_{\rm EGW}=\frac{0.01E}{|D^{\rm GW}_{E}|},~~\delta t_{\rm JGW}=\frac{0.005J}{|D^{\rm GW}_{J}|}$ (A6)
When the eccentricity is high, it is no longer necessary to decrease the time step as long as the pericenter passage is resolved. Therefore, we set a maximum of $0.2P$ on the time step in Equation A5 when $j\leq 0.4$ ($e\gtrsim 0.9$).
A.2 For orbits unbound to the MBH
$\begin{gathered}\delta E^{\rm GW}=D^{GW}_{E}P=-\frac{64}{5}\frac{M_{\bullet}^{2}m(m+M_{\bullet})^{1/2}G^{7/2}}{r_{p}^{7/2}c^{5}(1+e_{K})^{7/2}}\times\left[\left(\pi-\theta_{0}\right)\left(1+\frac{73}{24}e^{2}_{K}+\frac{37}{96}e^{4}_{K}\right)+\frac{(e^{2}_{K}-1)^{1/2}}{144}\left(301+\frac{673}{2}e^{2}_{K}\right)\right],\\ \delta J^{\rm GW}=D^{GW}_{J}P=-\frac{64}{5}\frac{M_{\bullet}^{2}mG^{3}}{r_{p}^{2}c^{5}(1+e_{K})^{2}}\times\left[(\pi-\theta_{0})\left(1+\frac{7}{8}e^{2}_{K}\right)+\frac{e_{K}}{8}\sin\theta_{0}(13+e^{2}_{K})\right],\end{gathered}$ (A7)
respectively, where $\cos\theta_{0}=1/e_{K}$. Similarly to Appendix A.1, here we damp the energy and angular momentum only if the particle has passed the pericenter of its orbit.
Appendix B Solving for the Relativistic ISO
Figure B: Left panel: The pericenter distance $r_{p}$ (in units of $r_{g}$) of an innermost stable orbit (ISO) in Boyer-Lindquist (BL) coordinates, plotted as a function of the orbital semi-major axis, defined as $a_{\rm GR}=(r_{p}+r_{a})/2$, where $r_{a}$ is the apocenter distance. Both $r_{p}$ and $r_{a}$ are solved using the method described in Appendix B. The colored lines show the results for the case of $a=1$ with various inclination angles ii. The solid black line shows the result for a Schwarzschild MBH. Middle panel: The mapped Newtonian pericenter distance $r_{\rm p,NW}$ of the ISO as a function of the Newtonian semi-major axis $a_{\rm NW}$, both defined by Equation C1. Right panel: The Newtonian pericenter distance $r_{\rm p,NW}$ of the ISO as a function of the Newtonian inclination angle $i_{\rm NW}$, calculated for orbits with a semi-major axis of $a_{\rm NW}=10^{4}r_{g}$.